You have a piece of property that measures 3/8 mile by 1/4 of a mile. What is the perimeter in yards?
1
Expert's answer
2012-10-08T10:40:09-0400
P=2(3/8+1/4)=2(3/8+2/8)=2∙5/8=5/4=1 1/4=1.25mile